# Runtime Safety through Adaptive Shielding: From Hidden Parameter Inference to Provable Guarantees

### v1

@@ Appendix B Proofs
We assume that the obstacle position $X_{t}$ can be derived from the state $s_{t}$. This assumption is reasonable, as the agent’s state typically includes safety-critical information. For instance, robots in navigation tasks use sensors to detect nearby obstacles, with this information integrated into the agent’s state. This setup applies to all navigation environments in Safety Gymnasium. Formally, we assume a 1-Lipschitz continuous function $\text{sensor}:S\to\mathbb{R}^{2M}$, defined as $\text{sensor}(s_{t})=[X_{1,t},X_{2,t},\ldots,X_{M,t}]$, where $X_{i,t}\in\mathbb{R}^{2}$ is the position of the $i$-th obstacle detected by the robot, and $M$ is the number of detected obstacles. When more than $M$ obstacles are present, the sensor typically detects the $M$ closest ones.
We adopt a Gaussian policy $\pi$, commonly employed in training RL policies across various algorithms.
Remark 1.
To ensure the validity of conformal prediction, we note that the parameter $\theta$, sampled at the start of each episode, remains fixed in each episode. This ensures that the dataset within each episode is exchangeable with respect to the transition dynamics $T_{\theta}$. For the calibration set, we collect online samples to maintain exchangeability. Specifically, during the first 100 steps, we gather samples for calibration without using the ACP region. After 100 steps, we employ the online-collected calibration set to determine the ACP region.
@@
Let $\hat{f}$ be a transition dynamics predictor and $h(a)=\min_{\hat{X}_{t+1}}\|\text{pos}(\hat{f}(s_{t},a))-\hat{X}_{t+1}\|$. Under the adaptive shielding mechanism with sampling size $N$ for each episode with parameter $\theta$, one of the following conditions holds:
@@
    $\mathbb{P}(s_{t+1}\in S_{\text{safe}}(s_{t},a_{t}))\geq 1-\delta$, where $s_{t+1}\sim T(\cdot\mid s_{t},a_{t},\theta)$,
@@
    $\min_{X_{t+1}}\|\text{pos}(s_{t+1})-X_{t+1}\|\geq\max_{a\in A}h(a)-\epsilon_{N}-2\Gamma_{t}$,
@@
Note that $s_{t+1}$ is safe if $\min_{X_{t+1}}\left\|\operatorname{pos}\left(s_{t+1}\right)-X_{t+1}\right\|>d.$ The ACP gives us a probabilistic bound on the deviation between the true next state $s_{t+1}$ and the predicted state $\hat{s}_{t+1}=\hat{f}(s_{t},a_{t})$ :
We connect the safety of $s_{t+1}$ to the position of the predicted state $\hat{s}_{t+1}$, using this bound. By triangle inequality, we have
$\min_{X_{t+1}}\left\|\operatorname{pos}\left(s_{t+1}\right)-X_{t+1}\right\|$ $\geq\min_{X_{t+1}}\|\operatorname{pos}\left(\hat{s}_{t+1}\right)-\hat{X}_{t+1}\|-\|\hat{X}_{t+1}-X_{t+1}\|-\left\|\operatorname{pos}\left(s_{t+1}\right)-\operatorname{pos}\left(\hat{s}_{t+1}\right)\right\|.$ (13)
@@
$\left\|\operatorname{pos}\left(s_{t+1}\right)-\operatorname{pos}\left(\hat{s}_{t+1}\right)\right\|\leq\left\|s_{t+1}-\hat{s}_{t+1}\right\|\text{~{} and ~{}}\|X_{t+1}-\hat{X}_{t+1}\|\leq\left\|s_{t+1}-\hat{s}_{t+1}\right\|.$
Hence, if $\left\|s_{t+1}-\hat{s}_{t+1}\right\|\leq\Gamma_{t}$ (which occurs with probability at least $1-\delta$ ), then
@@
$\hat{A}_{\text{safe}}\left(s_{t}\right)=\left\{a\in A\mid h(a)>d+2\Gamma_{t}\right\}.$
@@
Case 2: If we cannot select $a_{t}\in\hat{A}_{\text{safe}}(s_{t})$, our adaptive shielding mechanism samples $N$ actions $\{a^{(i)}_{t}\}$ and picks the action $a_{t}$ such $h(a_{t})=\max_{a\in\{a^{(i)}_{t}\}}h(a)$. Note that $h(a)$ is continuous on $a$ and a Gaussian policy $\pi$ assigns positive probability to any subset of action space $A$. Hence, as sample size $N$ goes to $\infty$, $\max_{a\in A}h(a)-h(a_{t})=\epsilon_{N}$ goes to 00. Also, by Equation 14, we have
$\|\text{pos}(s_{t+1})-X_{t+1}\|\geq h(a_{t})-2\Gamma_{t}=\max_{a\in A}h(a)-\epsilon_{N}-2\Gamma_{t}.$
@@
To prove the theorem, we recall the function $h(a)=\min_{\hat{X}_{t+1}}\|\text{pos}(\hat{f}(s_{t},a))-\hat{X}_{t+1}\|$, representing the minimum distance between the predicted state and predicted obstacles. Using this, we define the safe action set for the predicted state as:
$\hat{A}_{\text{safe}}\left(s_{t}\right)=\left\{a\in A\mid h(a)>d+2\Gamma_{t}\right\}.$
@@
Theorem 1. Given a Constrained Hidden Parameter MDP =$\mathcal{M}=\left\{S,A,\Theta,T,R,C,\gamma,P_{\Theta}\right\}$ with initial state $s_{0}\in S$, and failure probability $\delta\in(0,1)$, an optimal policy $\pi^{*}:S\times\Phi\rightarrow A$, augmented with an adaptive shield, maximizes the expected cumulative discounted return $J_{R}(\pi^{*})$, while satisfying the average cost rate constraint:
$\phi^{\pi^{*}}\left(s,\theta\right)=\lim_{H\rightarrow\infty}\frac{1}{H}\mathbb{E}_{\pi^{*},T_{\theta}}\left[\sum_{t=0}^{H-1}C(s_{t},a_{t},s_{t+1})\mid s_{0}=s,\theta\right]\leq\delta+\bar{\epsilon}(1-\delta),$ (15)
for some $0\leq\bar{\epsilon}\leq 1$ and $\theta\sim P_{\Theta}$.
@@
where $s_{t+1}\sim T(\cdot\mid s_{t},a_{t},\theta)$, and the safe state set is defined as:
$S_{\mathrm{safe}}\left(s_{t},a_{t}\right)=\left\{s^{\prime}\in S\mid C\left(s_{t},a_{t},s^{\prime}\right)=0,T\left(s^{\prime}\mid s_{t},a_{t},\theta\right)>0\right\}$
@@
$C\left(s_{t},a_{t},s_{t+1}\right)=\begin{cases}1&\text{ with probability at most }\delta\\ 0&\text{ with probability at least }1-\delta\end{cases},$
@@
By the linearity of expectation, this per-step bound extends to the long-term average cost for a fixed parameter $\theta$:
$\phi^{\pi^{*}}\left(s_{0},\theta\right)$ $=\lim_{H\rightarrow\infty}\frac{1}{H}\sum_{t=0}^{H-1}\mathbb{E}_{\pi^{*},T_{\theta}}\left[C\left(s_{t},a_{t},s_{t+1}\right)\right]$ (16)
@@
where $\bar{\epsilon}=\limsup_{H\rightarrow\infty}\frac{1}{H}\sum_{t=0}^{H-1}\mathbb{E}\left[\epsilon_{t}\right]$. This satisfies Equation 15, completing the proof. Moreover, if safe actions exist at each time step $t$, i.e., $\epsilon_{t}=\mathbb{P}(\hat{A}_{\text{safe}}=\emptyset)=0$, $\bar{\epsilon}$ becomes 00. Hence, we can bound the equation with a small failure probability $\delta$. ∎
Appendix C Additional Results for RQ1
In the main text, we presented results for the Car robot. Here, we provide results for the Point robot. Consistent with our earlier observations, Figure 3 shows that our method reduces safety violations with only a small reduction in reward, demonstrating robustness across different robotic platforms.
Figure 3: Training Curves for Safe Gymnasium (Point Robot, TRPO-Lag Base). (Top) Average episodic return versus training steps. (Bottom) Average episodic cost rate versus training steps.
Appendix D Adaptive Shielding with PPO-Lag
Figure 4: Training Curves for Safe Gymnasium (Car Robot, PPO-Lag Base). (Top) Average episodic return versus training steps. (Bottom) Average episodic cost rate versus training steps.
Figure 5: Training Curves for Safe Gymnasium (Point Robot, PPO-Lag Base). (Top) Average episodic return versus training steps. (Bottom) Average episodic cost rate versus training steps.
As our method is compatible with any RL policy $\pi$, we examine its impact when applied to a PPO-Lagrangian-based policy. Figures 4 and 5 demonstrate that integrating our method with the underlying policy consistently reduces safety violations.

### v2

@@ Appendix A Proofs
We assume that the obstacle position $X_{t}$ can be derived from the state $s_{t}$. This assumption is reasonable, as the agent’s state typically includes safety-critical information. For instance, robots in navigation tasks use sensors to detect nearby obstacles, with this information integrated into the agent’s state. This setup applies to all navigation environments in Safety Gymnasium. Formally, we assume a 1-Lipschitz continuous function $\text{sensor}:S\to\mathbb{R}^{2M}$, defined as $\text{sensor}(s_{t})=[X_{1,t},X_{2,t},\ldots,X_{M,t}]$, where $X_{i,t}\in\mathbb{R}^{2}$ is the position of the $i$-th obstacle detected by the robot, and $M$ is the number of detected obstacles. When more than $M$ obstacles are present, the sensor typically detects the $M$ closest ones. We adopt a Gaussian policy $\pi$, commonly employed in training RL policies across various algorithms.
Remark 1 (Validity under Sequential Dependence).
Standard conformal prediction assumes exchangeability of calibration and test data, an assumption that is generally violated in sequential decision-making due to temporal dependence. In our setting, a hidden context parameter $\phi$ is sampled at the beginning of each episode and remains fixed, so that the transition kernel $T_{\phi}$ is stationary within the episode. However, the marginal distribution of $(s_{t},a_{t})$ evolves over time as the policy interacts with the environment, and thus the resulting transition tuples are not exchangeable.
To account for this non-exchangeability, we use adaptive conformal prediction (ACP) (gibbs2021adaptive) as an online calibration mechanism. Our online calibration procedure is designed to control this nonstationarity by restricting calibration and test samples to the same episode and transition model. Specifically, during the first $100$ steps of each episode, we collect a calibration set of transition tuples without applying the ACP region. After this warm-up phase, we use the online-collected calibration data to compute the adaptive conformal threshold $\Gamma_{t}$ and construct the ACP region.
@@
Let $\hat{f}$ be a transition dynamics predictor and $e(a)=\min_{\hat{X}_{t+1}}\|\text{pos}(\hat{f}(s_{t},a))-\hat{X}_{t+1}\|$. Under the adaptive shielding mechanism with sampling size $N$ for each episode with parameter $\phi$, one of the following conditions holds for $s_{t+1}\sim T(\cdot\mid s_{t},a_{t},\phi)$:
@@
    $\mathbb{P}(s_{t+1}\in S_{\text{safe}}(s_{t},a_{t}))\geq 1-\delta$,
@@
    $\min_{X_{t+1}}\|\text{pos}(s_{t+1})-X_{t+1}\|\geq\max_{a\in A}e(a)-\epsilon_{N}-2\Gamma_{t}$,
@@
Note that $s_{t+1}$ is safe if
$\min_{X_{t+1}}\left\|\operatorname{pos}\left(s_{t+1}\right)-X_{t+1}\right\|>d.$
The ACP gives us a probabilistic bound on the deviation between the true next state $s_{t+1}$ and the predicted state $\hat{s}_{t+1}=\hat{f}(s_{t},a_{t})$ :
$\mathbb{P}\left(\|\hat{s}_{t+1}-s_{t+1}\|\leq\Gamma_{t}\right)\geq 1-\delta.$
Using this bound, we connect the safety of $s_{t+1}$ to the position of the predicted state $\hat{s}_{t+1}$. By triangle inequality, we have
$\min_{X_{t+1}}$ $\left\|\operatorname{pos}\left(s_{t+1}\right)-X_{t+1}\right\|\geq\min_{X_{t+1}}\|\operatorname{pos}\left(\hat{s}_{t+1}\right)-\hat{X}_{t+1}\|$
$-\|\hat{X}_{t+1}-X_{t+1}\|-\left\|\operatorname{pos}\left(s_{t+1}\right)-\operatorname{pos}\left(\hat{s}_{t+1}\right)\right\|.$ (13)
@@
$\left\|\operatorname{pos}\left(s_{t+1}\right)-\operatorname{pos}\left(\hat{s}_{t+1}\right)\right\|\leq\left\|s_{t+1}-\hat{s}_{t+1}\right\|\\$
and
$\|X_{t+1}-\hat{X}_{t+1}\|\leq\left\|s_{t+1}-\hat{s}_{t+1}\right\|.$
Hence, if $\left\|s_{t+1}-\hat{s}_{t+1}\right\|\leq\Gamma_{t}$ (which happens with probability at least $1-\delta$), then
@@
$\hat{A}_{\text{safe}}\left(s_{t}\right)=\left\{a\in A\mid e(a)>d+2\Gamma_{t}\right\}.$
@@
Case 2: If we cannot select $a_{t}\in\hat{A}_{\text{safe}}(s_{t})$, our adaptive shielding mechanism samples $N$ actions $\{a^{(i)}_{t}\}$ and picks the action $a_{t}$ such $e(a_{t})=\max_{a\in\{a^{(i)}_{t}\}}e(a)$. Note that $e(a)$ is continuous on $a$ and a Gaussian policy $\pi$ assigns positive probability to any subset of action space $A$. Hence, as sample size $N$ goes to $\infty$, $\max_{a\in A}e(a)-e(a_{t})=\epsilon_{N}$ goes to 0. Also, by Equation 14, we have
$\|\text{pos}(s_{t+1})-X_{t+1}\|\geq e(a_{t})-2\Gamma_{t}=\max_{a\in A}e(a)-\epsilon_{N}-2\Gamma_{t}.$
@@
To prove the theorem, we recall the function $e(a)=\min_{\hat{X}_{t+1}}\|\text{pos}(\hat{f}(s_{t},a))-\hat{X}_{t+1}\|$, representing the minimum distance between the predicted state and predicted obstacles. Using this, we define the safe action set for the predicted state as:
$\hat{A}_{\text{safe}}\left(s_{t}\right)=\left\{a\in A\mid e(a)>d+2\Gamma_{t}\right\}.$
@@
Theorem 1. Given a Constrained Hidden Parameter MDP $\mathcal{M}=(S,A,\Phi,T,R,C,\gamma,P_{\phi})$ with initial state $s_{0}\in S$, and failure probability $\delta\in(0,1)$, an optimal policy $\pi^{*}:S\times\Phi\rightarrow A$, augmented with an adaptive shield, maximizes the expected cumulative discounted return $J_{R}(\pi^{*})$, while satisfying the average cost rate constraint with an initial state $s$:
$\xi^{\pi^{*}}\left(s,\phi\right)$ $=\lim_{H\rightarrow\infty}\frac{1}{H}\mathbb{E}_{\pi^{*},T_{\phi}}\left[\sum_{t=0}^{H-1}C(s_{t},a_{t},s_{t+1})\mid s,\phi\right]$
$\leq\delta+\bar{\epsilon}(1-\delta),$ (15)
for some $0\leq\bar{\epsilon}\leq 1$ and $\phi\sim P_{\Phi}$.
@@
where $s_{t+1}\sim T(\cdot\mid s_{t},a_{t},\phi)$, and the safe state set is defined as:
$S_{\mathrm{safe}}\left(s_{t},a_{t}\right)=\left\{s^{\prime}\mid C\left(s_{t},a_{t},s^{\prime}\right)=0,T_{\phi}(s^{\prime}\mid s_{t},a_{t})>0\right\}$
@@
[…]
@@
By the linearity of expectation, this per-step bound extends to the long-term average cost for a fixed parameter $\phi$:
$\xi^{\pi^{*}}\left(s_{0},\phi\right)$ $=\lim_{H\rightarrow\infty}\frac{1}{H}\sum_{t=0}^{H-1}\mathbb{E}_{\pi^{*},T_{\phi}}\left[C\left(s_{t},a_{t},s_{t+1}\right)\right]$ (16)
@@
where $\bar{\epsilon}=\limsup_{H\rightarrow\infty}\frac{1}{H}\sum_{t=0}^{H-1}\mathbb{E}\left[\epsilon_{t}\right]$. This satisfies Equation A, completing the proof. Moreover, if safe actions exist at each time step $t$, i.e., $\epsilon_{t}=\mathbb{P}(\hat{A}_{\text{safe}}=\emptyset)=0$, $\bar{\epsilon}$ becomes 0. Hence, we can bound the equation with a small failure probability $\delta$. ∎
Appendix B Compatibility of Safety-Regularized Optimization with TRPO and CPO
This section provides a formal extension of the monotonic improvement guarantee of Trust Region Policy Optimization (TRPO) (schulman2015trpo) to our proposed safety-regularized objective, $J_{\text{aug}}(\pi)$. We first recap the foundational theorem of TRPO and then prove that this guarantee directly applies to our augmented objective. Finally, we provide a rigorous analysis of the trade-off between reward and safety that this guarantee implies
This section shows that incorporating SRO into TRPO and CPO preserves their standard monotonic improvement and worst-case constraint bounds. Our analysis follows directly from the original proofs and relies on interpreting SRO as a modification of the scalar advantage signal, rather than as a change to the underlying policy update or trust-region structure.
We formally define the augmented objective as $J_{R}(\pi)=\mathbb{E}\left[Q_{R}^{\pi}\right],J_{\text{safe}}(\pi)={\mathbb{E}}[Q^{\pi}_{\text{safe}}]$ and $J_{\text{aug }}(\pi)=\mathbb{E}[Q_{\text{aug }}^{\pi}]$. We restate the policy improvement guarantee of TRPO with SRO. During the optimization, KL constraint plays the same role as in standard TRPO. This only limits how far the updated policy is allowed to move from the current policy in one step, and is independent of whether we optimize the original reward $J_{R}$ or our safety-augmented objective $J_{\text{aug}}$.
Theorem 2 (schulman2015trpo). Let $L_{\pi}^{\text{aug }}(\tilde{\pi})=J_{\text{aug}}(\pi)+\mathbb{E}_{\phi\sim P_{\phi}s\sim\rho_{\pi},a\sim\tilde{\pi}}\left[A_{\text{aug }}^{\pi}(s,a,b_{\phi})\right]$. The performance of the new policy $\tilde{\pi}$ is lower-bounded by:
[…]
where $C_{\text{aug }}=\frac{2\gamma}{(1-\gamma)^{2}}\max_{s,a,\phi}\left|A_{\text{aug }}^{\pi}(s,a,b_{\phi})\right|$.
Monotonic improvement condition.. If the update $\tilde{\pi}$ satisfies
$L_{\pi}^{\operatorname{aug}}(\tilde{\pi})-C_{\text{aug }}D_{KL}^{\max}(\pi,\tilde{\pi})\geq J_{\text{aug }}(\pi),$
[…]
Interpretation of the Augmented Objective Trade-off. Let $\Delta J_{R}=J_{R}(\tilde{\pi})-J_{R}(\pi)$ and $\Delta J_{\text{safe}}=J_{\text{safe}}(\tilde{\pi})-J_{\text{safe}}(\pi)$. Under the monotonic-improvement condition above, for any valid policy update step:
$\Delta J_{R}+\alpha\cdot\Delta J_{\text{safe}}\geq 0$
This inequality provides a formal characterization of the trade-off between reward and safety.
Bounded Reward Degradation for Safety Improvement. If an update improves the safety objective ($\Delta J_{\text{safe}}>0$), we can rearrange the inequality to bound the permissible change in reward:
$\frac{\Delta J_{R}}{\Delta J_{\text{safe}}}\geq-\alpha\quad\Longrightarrow\quad\Delta J_{R}\geq-\alpha\cdot\Delta J_{\text{safe}}$
This proves that for a given gain in safety, the reward is guaranteed not to decrease by more than $\alpha$ times that gain. The hyperparameter $\alpha$ thus acts as a maximum acceptable cost in reward for a unit of safety improvement.
Bounded Safety Degradation for Reward Improvement. Conversely, if an update improves the reward objective ($\Delta J_{R}>0$), we can bound the permissible change in the safety term:
$\Delta J_{\text{safe}}\geq-\frac{1}{\alpha}\Delta J_{R}$
This proves that for a given gain in reward, the safety term is guaranteed not to decrease by more than $1/\alpha$ times that gain. This demonstrates that the algorithm will forgo policy updates that yield high rewards at the expense of excessive safety violations, where the threshold for excessive is explicitly controlled by $\alpha$.
Now, we also consider safety perspectives of SRO. We mainly analyze how SRO affects safety perspective considering theorems from CPO (achiam2017constrained).
Relation to CPO-style constraint guarantees. CPO’s worst-case bound (achiam2017constrained) depends only on KL constraint and cost advantage $A_{C}^{\pi_{k}}$. This is agnostic to how the updated policy $\pi_{k+1}$ is produced. Therefore, the bound from Proposition 2 in (achiam2017constrained)
$J_{C}\left(\pi_{k+1}\right)\leq J_{C}\left(\pi_{k}\right)+\frac{\sqrt{2\delta}\gamma}{(1-\gamma)^{2}}\epsilon_{C}^{\pi_{k+1}}$
remains the same with $J_{\text{aug}}$. The sole role of SRO is to restrict the updated policy $\pi_{k+1}$ so that the quantity $\epsilon_{C}^{\pi_{k+1}}=\max_{s}\left|\mathbb{E}_{a\sim\pi_{k+1}}[A_{C}^{\pi_{k}}(s,a)]\right|$ remains small. Because SRO penalizes actions with high estimated long-term cost, $\epsilon_{C}^{\pi_{k+1}}$ can be related to the safety weight $\alpha$ under the local approximation in (B). Substituting this relation into the CPO bound yields a worst-case cost bound that scales inversely with $\alpha$. Substituting this bound into the CPO inequality provided a worst-case cost guarantee for the SRO-TRPO update.
Throughout this section, we assume that the ball $B(a,\epsilon)$ is small enough that local averages approximate point values, i.e., $Q^{\pi}_{\text{safe}}(s,a)\approx-\frac{Q^{\pi}_{C}(s,a)}{V_{C}^{\pi}(s)+\epsilon}$. We begin with introducing necessary Lemma for bound of $A_{\text{aug}}^{\pi}$.
Lemma 2.
Let $A_{\mathrm{aug}}^{\pi}(s,a)=A_{R}^{\pi}(s,a)+\alpha A_{\mathrm{safe}}^{\pi}(s,a),$ where $A_{R}^{\pi}(s,a)=Q_{R}^{\pi}(s,a)-V_{R}^{\pi}(s),$
$A_{\text{safe }}^{\pi}(s,a)=Q_{\text{safe }}^{\pi}(s,a)-V_{\text{safe }}^{\pi}(s),$
and
$V_{\text{safe }}^{\pi}(s)=\mathbb{E}_{a^{\prime}\sim\pi(\cdot\mid s)}\left[Q_{\text{safe}}^{\pi}(s,a^{\prime}\right)].$
Then, $Q_{\text{safe }}^{\pi}(s,a)\in(-1,0]$ for all $(s,a)$, and $\left|A_{\mathrm{aug}}^{\pi}(s,a)\right|\leq\left|A_{R}^{\pi}(s,a)\right|+\alpha$
Proof.
By design, $Q_{\text{safe }}^{\pi}(s,a)\in(-1,0]$. Consequently, the value function $V_{\text{safe }}^{\pi}(s)$, being an expectation of $Q$, is also in ( 1,0-1,0 ]. The advantage is defined as $A_{\text{safe }}^{\pi}(s,a)=Q_{\text{safe }}^{\pi}(s,a)-V_{\text{safe }}^{\pi}(s)$. The maximum possible value is 0(1)=10-(-1)=1 (when Q=0,V=1Q=0,V=-1 ). The minimum possible value is 10=1-1-0=-1 (when Q=1,V=0Q=-1,V=0). Thus, $\left|A_{\text{safe}}^{\pi}(s,a)\right|\leq 1$. By triangular inequality, we have $\left|A_{\mathrm{aug}}^{\pi}(s,a)\right|=|A^{\pi}_{R}(s,a)+\alpha A_{\text{safe}}^{\pi}(s,a)|\leq\left|A_{R}^{\pi}(s,a)\right|+\alpha.$
We will split $A_{\text{aug}}^{\pi}$ into two parts $A^{\pi}_{R}$ and $A^{\pi}_{C}.$ Since the ball $B(a,\epsilon)$ is small enough that local averages approximate point values, our construction reduces to
$A_{\text{safe}}^{\pi_{k}}(s,a)$ $=Q_{\text{safe}}^{\pi_{k}}(s,a)-V_{\text{safe}}^{\pi_{k}}(s)\approx-\frac{Q_{C}^{\pi_{k}}(s,a)-V_{C}^{\pi_{k}}(s)}{V_{C}^{\pi_{k}}(s)+\varepsilon}$
$=-\frac{A_{C}^{\pi_{k}}(s,a)}{V_{C}^{\pi_{k}}(s)+\varepsilon}.$ (19)
Hence, we have
$A_{\text{aug}}^{\pi_{k}}(s,a)$ $=A_{R}^{\pi_{k}}(s,a)+\alpha A_{\mathrm{safe}}^{\pi_{k}}(s,a)$
$\approx A_{R}^{\pi_{k}}(s,a)-\lambda(s)A_{C}^{\pi_{k}}(s,a),$ (20)
where
$\lambda(s)=\frac{\alpha}{V_{C}^{\pi_{k}}(s)+\epsilon}.$ (21)
We assume that the reward and cost are bounded by $R_{\text{max}}$ and $C_{\text{max}}$, respectively. Then, for all $s,a$, we have
$|A_{R}^{\pi_{k}}(s,a)|\leq\frac{R_{\text{max}}}{1-\gamma},|A_{C}^{\pi_{k}}(s,a)|\leq\frac{C_{\text{max}}}{1-\gamma}$ (22)
$V_{C}^{\pi_{k}}(s)\leq\frac{C_{\text{max}}}{1-\gamma}.$ (23)
This implies that, for all $s$, we have
$\lambda(s)=\frac{\alpha}{V_{C}^{\pi_{k}}(s)+\epsilon}\geq\frac{\alpha}{\frac{C_{\max}}{1-\gamma}+\epsilon}=\frac{\alpha(1-\gamma)}{C_{\text{max}}+\epsilon(1-\gamma)}.$ (24)
Moreover, by rearranging Equation B, we have
$A_{C}^{\pi_{k}}(s,a)\approx\frac{A_{R}^{\pi_{k}}(s,a)-A_{\text{aug}}^{\pi_{k}}(s,a)}{\lambda(s)}.$
Combining Equations 22 and 24, and Lemma 2, we have
$|A_{C}^{\pi_{k}}(s,a)|$ $\leq\frac{1}{\lambda(s)}\left(|A_{R}^{\pi_{k}}(s,a)|+|A_{\text{aug}}^{\pi_{k}}(s,a)|\right)$ (25)
$\leq\frac{1}{\lambda(s)}\left(\frac{R_{\text{max}}}{1-\gamma}+\frac{R_{\text{max}}}{1-\gamma}+\alpha\right)$ (26)
$\leq\frac{C_{\text{max}}+\epsilon(1-\gamma)}{\alpha(1-\gamma)}\left(\frac{2R_{\text{max}}}{1-\gamma}+\alpha\right)$ (27)
$=\frac{2R_{\text{max}}(C_{\text{max}}+\epsilon(1-\gamma))}{\alpha(1-\gamma)^{2}}+\frac{C_{\text{max}}}{(1-\gamma)}+\epsilon.$ (28)
Thus, we have
$|A_{C}^{\pi_{k}}(s,a)|\leq O\left(\frac{1}{\alpha}\right)+\frac{C_{\text{max}}}{(1-\gamma)}+\epsilon.$ (29)
By the definition of $\epsilon_{C}^{\pi_{k+1}}=\max_{s}\left|\mathbb{E}_{a\sim\pi_{k+1}}[A_{C}^{\pi_{k}}(s,a)]\right|$, and Equation 29, CPO worst-case bound (Proposition 2 in (achiam2017constrained))
$J_{C}(\pi_{k+1})\leq J_{C}(\pi_{k})+\frac{\sqrt{2\delta}\gamma}{(1-\gamma)^{2}}\varepsilon_{C}^{\pi_{k+1}}$ (30)
becomes
$J_{C}(\pi_{k+1})\leq J_{C}(\pi_{k})+O\!\left(\frac{\sqrt{\delta}}{\alpha}\right)+\frac{C_{\text{max}}\sqrt{2\delta\gamma}}{(1-\gamma)^{3}}+\epsilon\frac{\sqrt{2\delta}\gamma}{(1-\gamma)^{2}},$ (31)
showing that stronger safety regularization (larger $\alpha$) tightens the worst-case constraint-violation bound for a fixed KL radius $\delta$.
Appendix C Pseudocode for the safety-regularized optimization for actor-critic
For clarity in this section only, we denote the parameters for policy, reward value critic, cost value critic, and cost Q-critic by $\theta$, $\phi$, $\psi$, and $\omega$, respectively. We use $sg(\cdot)$ to indicate a stop-gradient operation.
1:Initialize: Policy $\pi_{\theta}$, Reward Critic $V^{\phi}_{R}$, Cost Value Critic $V^{\psi}_{C}$, Cost Q-Critic $Q^{\omega}_{C}$.
2:Hyperparameters: Safety bonus coefficient $\alpha$, KL constraint $\delta_{KL}$.
3:for each training epoch do
4:  Collect trajectories $\mathcal{D}$ using current policy $\pi_{\theta}$.
5:  Compute targets $V_{R}^{\text{targ}},V_{C}^{\text{targ}}$ and advantages $A_{R},A_{C}$ using GAE.
6:  Store observations, actions, log-probs, and targets in buffer.
7:  for $k=1$ to $K$ update iterations do
8:   Sample mini-batch $\mathcal{B}=\{(s,a,\log\pi_{old},V^{\text{targ}},A)\}$ from $\mathcal{D}$.
9:   // 1. Update Reward Critic
10:   Minimize $\mathcal{L}_{V_{R}}=\mathbb{E}_{\mathcal{B}}\left[(V^{\phi}_{R}(s)-V_{R}^{\text{targ}})^{2}\right]$.
11:   // 2. Update Cost Critics ($V_{C}$ and $Q_{C}$)
12:   Compute $V_{stop\_grad}=sg(V^{\psi}_{C}(s))$.
13:   Minimize $\mathcal{L}_{V_{C}}=\mathbb{E}_{\mathcal{B}}\left[(V_{stop\_grad}-V_{C}^{\text{targ}})^{2}\right]$.
14:   Minimize $\mathcal{L}_{Q_{C}}=\mathbb{E}_{\mathcal{B}}\left[((Q^{\omega}_{C}(s,a)-V_{stop\_grad})-A_{C})^{2}\right]$.
15:   // 3. Compute Safety Estimates
16:   Sample NN noise vectors $\epsilon\sim\mathcal{N}(0,\sigma)$ to generate local actions $a^{\prime}=a+\epsilon$.
17:   Estimate local Cost Q-values $Q_{approx}=\frac{1}{N}\sum_{a^{\prime}}\pi(a^{\prime}\mid s)Q^{\omega}_{C}(s,a^{\prime})$.
[…]
19:   // 4. Update Actor
20:   Compute surrogate advantage:
$A_{\text{aug}}=A_{R}+\alpha\cdot Q_{\text{safe}}$
21:   Update $\theta$ by maximizing policy objective $\mathcal{L}_{\pi}$ using $A_{\text{aug}}$.
22:  end for
23:end for
Algorithm 1 Actor-Critic with Safety-Regularized Objective
Appendix D Connection to Control Theory
In this section, we introduce a brief overview of barrier certificate approaches in control theory and then relate these ideas to our shielding mechanism.
Recent advances in safe control and safe reinforcement learning suggest using Control Barrier Functions (CBFs) or barrier-like certificates to establish forward invariance of a safe set. Classical and neural CBF methods construct a differentiable barrier function $h(s)$ whose evolution satisfies a discrete- or continuous-time invariance condition, and use this certificate to guarantee safety under learned or partially known dynamics. This paradigm has been widely applied to safe RL, including disturbance-observer-based barrier methods (cheng2023safe), reachability-based approximations (ganai2023iterative), soft-barrier formulations for stochastic environments (wang2023enforcing), neural CBFs integrated directly into RL (xiao2023bnet), and the joint use of CBFs and control Lyapunov functions (CLFs) for enhanced stability and safety under model uncertainty (choi2020reinforcement). These approaches learn or optimize a barrier function jointly with the policy or the dynamics model, and safety depends on the existence of a valid barrier certificate, which is often difficult to find.
Control Barrier Function
A function $\alpha:[0,a)\to[0,\infty)$ is called a class-$\mathcal{K}$ function if it is continuous, strictly increasing, and satisfies $\alpha(0)=0$. In classical nonlinear control, forward invariance of a safe set $\mathcal{S}_{\text{safe}}\subseteq\mathcal{S}$ is certified through a Control Barrier Function (CBF) $h:\mathcal{S}\rightarrow\mathbb{R}$ that satisfies the differential constraint
$\dot{h}(s)+\alpha(h(s))\geq 0,$ (32)
for an extended class-$\mathcal{K}$ function $\alpha$. If (32) holds for all admissible controls, then trajectories starting in $\mathcal{S}_{\text{safe}}=\{s\mid h(s)\geq 0\}$ remain in that set for all future times, guaranteeing forward invariance. The shielding guarantee proposed in Theorem 1 is closely related to the notion of forward invariance in control theory.
Relation to our Shield.
Our adaptive shield provides an analogous guarantee in the discrete-time and data-driven setting. Rather than enforcing the differential condition (32), we bound the change in a safety function $\nu(e(s_{t}),E_{t})$ between successive steps using its Lipschitz continuity,
$\nu(e(s_{t+1}),E_{t+1})$ $\geq\nu(e(s_{t}),E_{t})-L_{\nu}\,\|e(s_{t+1})-e(s_{t})\|$
$\geq\nu(e(s_{t}),E_{t})-L_{\nu}\Delta_{\max}.$ (33)
Thus, whenever $\nu(e(s_{t}),E_{t})>L_{\nu}\Delta_{\max}$, the next state $s_{t+1}$ remains within the safe region. This is a discrete-time forward-invariance condition derived from the structure of the cost function $C(s_{t},a_{t},s_{t+1})$ and the Lipschitz property of $\nu$. The adaptive conformal bound $\Gamma_{t}$ introduced in the shield plays the role of a stochastic disturbance margin, producing a probabilistic forward invariance guarantee.
While neural and classical CBF methods construct a barrier certificate $h(s)$ satisfying the invariance condition Equation 32 either through analytical dynamics or by learning $h$ jointly with a dynamics model, our approach leverage the structure of the cost function such as Lipschitz property, and conformal prediction bound without learning a barrier function.